Shift Summary for Eve, 1/10/99


Subject: Shift Summary for Eve, 1/10/99
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:16:54 CST


              Shift Summary for eve shift, 10/Jan/99

On Shift: Bob Tschirhart, Angela Bellavance

Hours of beam delivered at all intensities: 8.2
Hours of beam delivered at nominal intensity: 8.2
Hours of E799 data (special runs not included): 4.6
Hours of special runs (laser scans, muon runs...): 0.5

Accounting of lost time: Tape problems, DAQ dead time, setup for special
                         runs, lost permit.

========================================================================
Runs ended this shift:
Run 15466 - 4M events - BEAM_799
Run 15467 - ? events - BEAM_MUON
Run 15472 - 494K events - BEAM_MUON
Run 15473 - PED
Run 15474 - 376K events - KAMI_KE3
Run 15475 - ongoing - KAMI_KE3

Tapes brought to FCC this shift: UPA313-UPA321
                                 UPB313-UPB321
                                 UPC313-UPC321
                                 UPS105-7
                                 UPE021
                                 UPF076-7
                                 UPH056
                                 UPK040
                                 UPM056
                                 UPMU07

========================================================================
Accesses during shift:

E. Blucher to work on regenerator.
B. Tschirhart, Rick Ford into NM2 to set up for Ke3 spec. run (KAMI_KE3).
B. Tschirhart, E. Blucher, M. Arenton into NM4 to set up for KAMI_KE3 run.
B. Tschirhart, R. Kessler to find lost permit.

========================================================================
Overview of shift:
  E799 datataking, prepare for Ke3 special run (KAMI_KE3), KAMI_KE3
datataking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
              15:30 Trying to get E799_BEAM running. Having problems
                        with tapes.
15466 1 15:45 Run started.
      18 16:08 Beam aborts for 1 spill.
      20 Laser repeater cycled. Missed 2 spills.
      22 Ktevana errors. KFREND_PLAY and QUARTZ RESET.
      26-8 16:19 Beam abort at F6. 3 spills lost.
      39 Beam gone for 1 spill.
              16:51 Tape UPX020 mounted on ktev3.
      69 17:16 Triggers hang due to memory being full. They
                        restart themselves a little way into the
                        spill. This occurs frequently until the end of
                        this run.
              17:18 Start getting CsI errors. They are guessed to be
                        readout errors and ignored.
              17:58 End run.
15467 1 18:05 Start muon run.
      3 End muon run.
              18:20 Start getting set up for special Ke3 run. MCR
                        called to steer beam away. Various people paged
                        to come in.
              19:01 Leo notices an ETOT problem. Jason paged -
                        looking into it.
15472 1 20:13 Muon run with tagged ke3 configuration started.
              20:27 Muon run ended.
15473 1 20:34 Pedistal run done.
15474 1 20:51 Start taking KAMI_KE3 data!
              21:11 Run aborts with a tape drive failure. Clean
                        drive, replace tapes.
15475 1 21:38 Start another KAMI_KE3 data run.
      61-2 Triggers hang. Laser repeater cycle + KFREND
      70 DYC full - KFREND
      76-96 23:30 Triggers hang again. KFREND and laser voodoo
                        don't work. Permit lost downstairs. Call MCR to
                        request access. HA latch and TDC crate was
                        holding the permit - bypassed.
      97 Beam restored.
      104 NM3S magnet restored (found it had tripped off).
              
========================================================================
Follow up Notes on Previous Shift:

no need to do "QUARTZ RESET"s anymore

========================================================================
Items to Follow up on Next Shift:

SDAQ has a red CAMAC button. The error reads: "Crate 16, rack 191, DTM
front panel switch is on 'tracking'". Actually, this module is completely
out of the readout, so ignore this error message.

Check out HA latch/TDC crate during next downtime. Repair and put back in
readout?

_______
Angela Bellavance work: (630) 840-2801
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
                                - Arthur C. Clarke



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:16:56 CST